CRY+VIS Interface Meeting
Date: 2022/9/30 13:00-13:25
Participants: Takahashi, Aso, Yokozawa, Ikeda, Mitsuhashi, Ushiba, Akutsu, Sato, Node
- Type-A: Offload of the BF GAS in IY. Replacement of the HPCD in EX. Test of the Observation mode.
- Type-B: Windshield in BS.
- Type-Bp: Windshield in PR2/PR3. Monitoring of the PRM in the next week.
- Type-C: Health check of the IMMT suspensions. Recovery of the OMMT1/2/OSTM suspensions. Final check of the OMMT/OSTM suspensions in the next week.
- Replacement of cables: EXT remains.
- TM height setting:
- Oplev: Windshields.
- ACC: Evaluation.
- Commissioning in vacuum: Locking for a long time.
Acceptance check: DropboxPaper
- Vacuum system: Closing of the IMMT-IFI-PRM chambers. Leak check.
- Heater control: Test in IY will be planned.
- TMS: Implementation of the viewport in TMSY in the next week.
Date: 2022/9/30 14:00-14:20
Participants: Takahashi, Washimi, Ushiba, Tanaka, Fabian, Aso
- Offloaded the BF GAS in IY.
- Evaluated the accelerometers for the IPs. The performance was worse than the prediction.
- We build a new windshield for the BS oplev.
- In SR2, the seismometer on ground was calibrated with respect to the IP LVDTs and geophones.
- In SR2, the sensor correction filter Terrence designed for SRM was copied to SRM filter banks for temporary usage.
- In BS, a leak was found in an electrical feedthrough, we disconnected the cables to facilitate this work.
- Took photos of the EQ stoppers and the damping magnets in the IMMT suspensions.
- Recovered the touching points in the OMMT/OSTM suspensions.
- Type-Bp: Monitoring of the PRM.
- Type-C: Final check of the OMMT/OSTM suspensions.
- Regulation to use the clean booth will be changed.
(Date in Kamioka)
The week from 10/3
- Takahashi: 10/3-7
- Hirata: 10/4-7
- Aso: 10/3-6
The week from 10/10